OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). y = 7x - 3 y = -1/7 x + 2 y = -7x + 10 y = 7x - 10

OpenStudy (danjs):

Can you tell what the slope of the given line is y = 7x + 2 ?

OpenStudy (danjs):

Slope intercept form of a line y = m*x+b m= slope and b= y-intercept

OpenStudy (danjs):

The slope of the give line ; y = 7x + 2; is 7. A line parallel to that given line will have the SAME slope. Slope of the parallel line also = 7

OpenStudy (anonymous):

i dont really know how to do any of it

OpenStudy (danjs):

Have point (x1,y1) = (1,-3) and slope m = 7 Use Point-Slope form for a line to get line y - y1 = m(x-x1) plug in the values of m, x1, y1

OpenStudy (anonymous):

so (x1,y1) = (1,-3) and slope m = 7?? idk

OpenStudy (danjs):

yes, just put those numbers into y - y1 = m*(x - x1)

OpenStudy (danjs):

Summary. the slope of the line given is 7 , it is in the form y=m*x + b, m=7. A parallel line will have the same slope.

OpenStudy (danjs):

y + 3 = 7(x - 1) y + 3 = 7x - 7 y = 7x - 10

OpenStudy (anonymous):

1-(-3)=m*7x + 2

OpenStudy (danjs):

y = 7x - 10 is parallel to y = 7x + 2

OpenStudy (danjs):

m = 7, you have to replace m with 7, you still had m in your equation

OpenStudy (texaschic101):

first you find the slope of the given equation. It is 7. now use y = mx + b form slope(m) = 7 (1,-3)...x = 1 and y = -3 now we sub and find b (the y intercept) -3 = 7(1) + b -3 = 7 + b -3 - 7 = b -10 = b so your parallel equation is : y = 7x - 10

OpenStudy (anonymous):

ok thank you i still dont really get it but thanks

OpenStudy (danjs):

A slope is how 'steep' the ramp of the line is. A parallel ramp will have the same 'steepness'